The density of a 3 M sodium thiosulphate solution (Na2S2O3) is 1.25 g per mL. Calculate (i) the percentage by weight of sodium thiosulphate (ii) the mole fraction of sodium thiosulphate and (iii) the molalities of Na + and S2O_3^(2-) ions.

Grade:10

1 Answers

Aditi Chauhan
9 years ago
Sol. Let us consider 1.0 L solution for all the calculation. (i) Weight of 1 L solution = 1250 g Weight of Na2S2O3 = 3 × 158 = 474 g ⇒ Weight percentage of Na2S2O3 = 474/1250 × 100 = 37.92 (ii) Weight of H2O in 1 L solution = 1250 – 474 = 776 g Mole-fraction = 3/(3+ 776/18) = 0.065 (iii) Molality of Na+ = (3 ×2)/776 × 100 = 7.73 m Molality of S2O_3^(2-) = (3 )/(776 ) × 1000 = 3.86 m